OpenStudy (anonymous):

For f(x)=2x-6, find f(2). A. -2 B. 2 C. -10 D. 10

Vocaloid (vocaloid):

replace "x" with "2" and find f(x)

OpenStudy (anonymous):

as equated= f(x) = (2)(2)-6 2x2 = 4 4 - 6 = -2

OpenStudy (anonymous):

so -2